Present value of tax savings from depreciation of robot in 10 years

Rate of return 12%

Income tax 30%

Cost of robot 40000

Disposal value of robot 1000

years 10

Straight line depreciation

How to do it on excel
Answer #1

I will give you the answer first, then the excel formula version

Year Depreciation Tax on Depreciation PV Factor Present Value
1 3900 1170 0.892857143 $      1,044.64
2 3900 1170 0.797193878 $         932.72
3 3900 1170 0.711780248 $         832.78
4 3900 1170 0.635518078 $         743.56
5 3900 1170 0.567426856 $         663.89
6 3900 1170 0.506631121 $         592.76
7 3900 1170 0.452349215 $         529.25
8 3900 1170 0.403883228 $         472.54
9 3900 1170 0.360610025 $         421.91
10 3900 1170 0.321973237 $         376.71
$ 39,000.00 $                11,700.00 $      6,610.76
Year Depreciation Tax on Depreciation PV Factor Present Value
1 =(40000-1000)/10 =D4*30% =1/(1+12%)^C4 =E4*F4
=C4+1 =(40000-1000)/10 =D5*30% =1/(1+12%)^C5 =E5*F5
=C5+1 =(40000-1000)/10 =D6*30% =1/(1+12%)^C6 =E6*F6
=C6+1 =(40000-1000)/10 =D7*30% =1/(1+12%)^C7 =E7*F7
=C7+1 =(40000-1000)/10 =D8*30% =1/(1+12%)^C8 =E8*F8
=C8+1 =(40000-1000)/10 =D9*30% =1/(1+12%)^C9 =E9*F9
=C9+1 =(40000-1000)/10 =D10*30% =1/(1+12%)^C10 =E10*F10
=C10+1 =(40000-1000)/10 =D11*30% =1/(1+12%)^C11 =E11*F11
=C11+1 =(40000-1000)/10 =D12*30% =1/(1+12%)^C12 =E12*F12
=C12+1 =(40000-1000)/10 =D13*30% =1/(1+12%)^C13 =E13*F13
=SUM(D4:D14) =SUM(E4:E14) =SUM(G4:G14)